40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Jan 13 03:00:01 2011, Jenne, Update, IOO, WFS shenanigans D990249-B-1_MCWFS_schematic.pdf
    Reply  Thu Jan 13 12:56:57 2011, rana, Update, IOO, WFS shenanigans PerkinElmerQPDs.pdf
    Reply  Fri Jan 28 13:07:31 2011, Jenne, Update, IOO, Beam is back on the WFS 
    Reply  Fri Jan 28 18:19:21 2011, Jenne, Update, IOO, WFS2 has some kind of oil on it Oily_WFS2.jpg
    Reply  Mon Feb 14 15:59:49 2011, Jenne, Update, IOO, WFS quantum efficiency as a function of angle 
       Reply  Wed Jun 15 00:49:34 2011, Suresh, Update, IOO, WFS2 has been fixed. WFS_QE_measurement.pngSensorsBrochure-p12.pdfP6150121.JPGP6150124.JPG
Message ID: 4819     Entry time: Wed Jun 15 00:49:34 2011     In reply to: 4289
Author: Suresh 
Type: Update 
Category: IOO 
Subject: WFS2 has been fixed. 

 

The WFS2 sensor head had a damaged Quadrant PIN diode (YAG-444-4A).  This has been replaced by a   YAG-444-4AH  which has a responsivity of 0.5 A/W. 

P6150121.JPG     P6150124.JPG

The responsivity of each quadrant was measured at normal incidence.  A diagram of the set up with the relevant power levels is attached.  The precision of these measurement is about 5% .  Largely because the power levels measured are sensitive to the position of the laser beam on the power meter sensor head (Ophir with ND filter mask taken off).  Putting the mask back on did not solve this problem.

The incident power was 0.491mW  of which about 0.026mW was reflected from the face of the QPD.  The beam was repositioned on the QPD to measure the response of each quadrant.  In each case the beam was positioned to obtain maximum DC output voltage from the relevant quadrant.  A small amount of spill over was seen in the other quadrants.  The measurements are given below

WFS2 DC output measurements (mV)
  Position 1 Position 2 Position 3 Position 4 Dark
Q1 244 6.7 5.4 6.9 4
Q2 5.9 238 8.4 5 5
Q3 9 6.6 236 7.3 6
Q4 7.5 7 7 252 7

WFS_QE_measurement.png

To measure these DC outputs of from the sensor-head a breakout board for the 25-pin D-type connector was used as in the previous measurements.  The results are given below

 

WFS2 Quantum Efficiency measurement

  DC out (mV)

Responsivity

A/W

Quantum Efficiency (%)
Q1 238 0.52 0.60
Q2 233 0.50 0.59
Q3 230 0.50 0.58
Q4 244 0.53 0.61

 

The measured responsivity agrees with the specification from the manufacturer.  It is to be noted that the previous QPD is reported to have a slightly smaller responsivity 0.4 A/W at 1064 nm.  The data sheet is attached. 

Since the new QPD may have a slightly different capacitance the RF transfer function of the WFS2 needs to be examined to verify the location of the resonances. 

 

Quote:

[Larisa and Jenne]

A few weeks ago (on the 28th of January) I had tried to measure the quantum efficiency of one quadrant of the WFS as a function of angle.  However, Rana pointed out that I was a spaz, and had forgotten to put a lens in front of the laser.  Why I forgot when doing the measurement as a function of angle, but I had remembered while doing it at normal incidence for all of the quadrants, who knows?

Anyhow, Larisa measured the quantum efficiency today.  She used WFS2, quadrant 1 (totally oil-free), since that was easier than WFS1.  She also used the Jenne Laser (with a lens), since it's more stable and less crappy than the CrystaLasers.  We put a 50 Ohm terminator on the RF input of the Jenne Laser, since we weren't doing a swept sine measurement.  Again, the Ophir power meter was used to measure the power incident on the diode, and the reflected power, and the difference between them was used as the power absorbed by the diode for the quantum efficiency measurement.  A voltmeter was used to measure the output of the diode, and then converted to current as in the quote below. 

Still on the to-do list:  Replace the WFS2 diode.  See if we have one around, otherwise order one.  Align beams onto WFS so we can turn on the servo.

QE = (h*c)/(lambda*e) * (I/P)

Where I = (Volts from Pin1 to GND)/2 /500ohms
P = Power from laser - power reflected from diode.
h, c, e are the natural constants, and lambda is 1064nm.
Also, I/P = Responsivity


Larissa is going to put her data and plots into the elog shortly....

Quote:

Quantum Efficiency Measurement:

I refer to Jamie's LHO elog for the equation governing quantum efficiency of photodiodes: LHO 2 Sept 2009

The information I gathered for each quadrant of each WFS was: [1] Power of light incident on PD (measured with the Ophir power meter), [2] Power of light reflected off the PD (since this light doesn't get absorbed, it's not part of the QE), and [3] the photo current output by the PD (To get this, I measured the voltage out of the DC path that is meant to go to EPICS, and backed out what the current is, based on the schematic, attached). 

I found a nifty 25 pin Dsub breakout board, that you can put in like a cable extension, and you can use clip doodles to look at any of the pins on the cable.  Since this was a PD activity, and I didn't want to die from the 100V bias, I covered all of the pins I wasn't going to use with electrical tape.  After turning down the 100V Kepco that supplies the WFS bias, I stuck the breakout board in the WFS.  Since I was able to measure the voltage at the output of the DC path, if you look at the schematic, I needed to divide this by 2 (to undo the 2nd op amp's gain of 2), and then convert to current using the 499 Ohm resistor, R66 in the 1st DC path.  

I did all 4 quadrants of WFS1 using a 532nm laser pointer, just to make sure that I had my measurement procedure under control, since silicon PDs are nice and sensitive to green.  I got an average QE of ~65% for green, which is not too far off the spec of 70% that Suresh found.

I then did all 8 WFS quadrants using the 1064nm CrystaLaser #2, and got an average QE of ~62% for 1064 (58% if I exclude 2 of the quadrants....see below).  Statistics, and whatever else is needed can wait for tomorrow.

Problem with 2 quadrants of WFS2?

While doing all of this, I noticed that quadrants 3 and 4 of WFS2 seem to be different than all the rest.  You can see this on the MEDM screens in that all 6 other quadrants, when there is no light, read about -0.2, whereas the 2 funny quadrants read positive values.  This might be okay, because they both respond to light, in some kind of proportion to the amount of light on them.  I ended up getting QE of ~72% for both of these quadrants, which doesn't make a whole lot of sense since the spec for green is 70%, and silicon is supposed to be less good for infrared than green.  Anyhow, we'll have to meditate on this.  We should also see if we have a trend, to check how long they have been funny.

 

 

Attachment 2: SensorsBrochure-p12.pdf  48 kB  | Hide | Hide all
SensorsBrochure-p12.pdf
ELOG V3.1.3-